# A Unified Industrial Large Knowledge Model Framework in Smart Manufacturing

### v1

@@ Abstract
The recent emergence of large language models (LLMs) shows the potential for artificial general intelligence, revealing new opportunities in industry 4.0 and smart manufacturing. However, a notable gap exists in applying these LLMs in industry, primarily due to their training on general knowledge rather than domain-specific knowledge. Such specialized domain knowledge is vital for effectively addressing the complex needs of industrial applications. To bridge this gap, this paper proposes an Industrial Large Knowledge Model (ILKM) framework emphasizing their potential to revolutionize the industry in smart manufacturing. In addition, ILKMs and LLMs are compared from eight perspectives. Finally, “6S Principle” is proposed as the guideline for the development of ILKMs in smart manufacturing.
@@ 1 Introduction
In the era of Industry 4.0 [5], a paradigm shift is unfolding in the manufacturing sector, driven by the advent of smart manufacturing practices. At the heart of this revolution lies the integration of many advanced techniques, such as Industrial Big Data Analytics [15], Industrial Artificial Intelligence (AI) [7], Machine Learning (ML) [11, 17], Cyber-Physical Systems (CPS) [6], Internet of Things (IoT) [13], etc. These techniques are all aimed at enhancing efficiency, productivity, and flexibility in manufacturing processes.
Recent advances in large language models (LLMs) [2, 14, 20] have showcased extraordinary capabilities in natural language processing for understanding, interpreting, and generating human language. However, a gap exists in the application of these LLMs within smart manufacturing, primarily because LLMs are predominantly trained on general knowledge, not domain knowledge, which may not be entirely suitable for the specific and often intricate needs of industrial applications. Therefore, there is an urgent need for the development of an advanced large ML model leveraging the powers of LLMs and domain knowledge to address complex challenges in smart manufacturing.
Recognizing this gap, we propose an industrial large knowledge model (ILKM) framework for domain-driven data-centric industrial systems in Smart Manufacturing, and the “6S Principle” is proposed for the development of ILKM. The role of ILKMs and the comparison with LLMs are discussed in detail. Through this exploration, we aim to provide a comprehensive understanding of the transformative power of ILKMs in the modern manufacturing landscape, highlighting their significance as a cornerstone of the ongoing industrial revolution.
@@ 2 The role of industrial large knowledge models in smart manufacturing
In the domain of smart manufacturing, the deployment of ILKMs emerges as a pivotal element. Figure 1 shows the general process of how ILKM works within Industry 4.0, where ILKM functions at the core of this advanced manufacturing paradigm. The process begins with the acquisition and aggregation of a vast array of industrial data, derived from diverse industrial products. This data is categorized into two primary forms: human-interpretable data and structured machine-generated data. Leveraging technologies such as LLMs, a comprehensive Large Knowledge Library (LKL), along with various Machine Learning (ML) and Artificial Intelligence (AI) techniques, the ILKM serves as Artificial General Intelligence (AGI), which plays a vital role in enabling advanced and sophisticated data analytics. This, therefore, paves the way for a more insightful and informed decision-making process. Beyond this, the ILKM can also interface with and improve supply chain management, leading to more efficient, resilient, and customer-focused operations. Additionally, the solutions generated by the ILKM undergo evaluation by Subject Matter Experts (SMEs), who play a crucial role in validating and refining the industrial data, thereby aiding in the continual optimization of ILKM’s output. This iterative process of assessment and feedback is integral to ensuring the relevance and effectiveness of ILKM’s solutions. Overall, the ILKM underscores the transformative potential of data-driven insights, offering detailed and comprehensive optimization, and enhancement directions for industrial products in smart manufacturing.
Figure 2: Industrial Large Knowledge Model Framework.
3 Comparison between industrial large knowledge models and large language models
The main difference between the ILKM and LLMs lies in their purpose and functionality. The ILKM is designed to handle specific industrial tasks, often integrating vast amounts of structured data and domain-specific knowledge to provide precise, expert-level insights, while LLMs are more generalized, leveraging extensive training on diverse textual data to solve language-related tasks like text generation, conversation, language translation, etc. To better illustrate the idea of ILKMs, a detailed comparison between ILKM and LLM is presented in Table 1. We compare and explain them from eight perspectives: data, purpose, applications, data privacy and security, domain-specific knowledge, integration and customization, scalability, and real-time decision-making.
@@ 3 Comparison between industrial large knowledge models and large language models
4 Industrial large knowledge model framework
The proposed ILKM framework, shown in Figure 2, provides a step-by-step guideline for developing and deploying ILKMs using industrial data for enhancing smart manufacturing capabilities in areas such as predictive maintenance, process monitoring, quality control, QA platform, data analytics, etc. The framework of the ILKM consists of four pivotal steps: (1) the construction of a LKL categorized by human-interpretable and structured machine-generated data; (2) the preparation of domain-specific instruction data; (3) the development of domain knowledge large language model based on the domain-specific data and domain instruction data; (4) the establishment of an intelligent multi-expert machine learning system. As illustrated in Figure 2, the details about the ILKM framework are outlined as follows:
4.1 Large knowledge library construction
The initial step of constructing an ILKM involves the creation of an LKL. This library is pivotal for accommodating the breadth and diversity of industrial data, thus serving as a foundational resource for subsequent analytical tasks. During this phase, it’s essential to systematically categorize the data into domain-specific categories. Such organization enables researchers and data scientists to streamline their efforts, allowing for efficient retrieval of domain-specific data to inform the development of machine learning models tailored to address distinct industry-related challenges. Within these categories, based on the usage and nature of the data, industrial data can be divided into two primary types: human-interpretable and structured machine-generated. On the one hand, human-interpretable data comprises elements such as textual documents, annotated images, coding scripts crafted by programmers, and multimedia content that are inherently designed for human cognition. These data can be seen as insightful information or knowledge and it can be used for later domain knowledge machine learning model development. On the other hand, structured machine-generated data, comprising sensor readings, machine logs, operational parameters, etc, is often leveraged for analytical purposes in technical and industrial contexts.
4.2 Domain instruction data preparation
In the second step of the ILKM framework, the focus shifts to transforming domain-specific data (human-interpretable data from LKL) into structured domain instruction sets by achieving multi-modal data fusion and generating domain-centric knowledge. This transformation is crucial for enhancing the performance of LLMs in targeted domains. These structured instructions, vital for the fine-tuning process of the LLM, ensure it is adept at addressing domain-specific challenges and improving problem-solving capabilities. As depicted in Figure 2, the domain instruction data is organized into three parts: firstly, the domain instruction, which identifies the problem’s domain and may include sub-tags for refined categorization; secondly, the input, which clearly outlines the current problem; and thirdly, the output, which presents the corresponding solution [2].
4.3 Domain knowledge large language model development
The third step of the ILKM framework entails an initial pre-training of the base LLM with domain-specific data sourced from LKL. This pre-training imbues the LLM with rich domain-specific insights. Following this, the pre-trained LLM undergoes a meticulous fine-tuning process, guided by domain instructions, to evolve into a Domain Knowledge LLM adept in the particularities of its designated field. To refine the LLM’s expertise, some common techniques to enhance and train LLMs can be summarized as follows: reinforcement learning from human feedback (RLHF) [4, 21], instruction tuning [10, 18], mixture of experts [8, 12], prompt engineering [1, 9, 19], and leveraging attention mechanisms [16], etc. The objective of this step is to build a robust LLM that possesses extensive domain knowledge and expertise and it could be used to guide the development of new machine-learning models that are capable of addressing complex challenges and real industrial problems.
4.4 Intelligent domain expert machine learning system
Upon the successful training of the Domain Knowledge LLM, the fourth step utilizes it as a domain expert for subsequent specialized model development. In this step, domain instruction data serves as the prompt, propelling the LLM to address specific analytical problems. The Domain Knowledge LLM, acting on these instruction inputs, proposes targeted solutions. Additionally, human experts may interact and intervene, offering strategic guidance to refine the LLM’s outputs. These solutions are then transferred to a coding-focused LLM [3], which incrementally develops code aligned with the Domain Knowledge LLM’s insights, thereby creating a new ML model for specific problems. In addition, the structured machine-generated data serves as the dataset for new ML model training and testing. Finally, this step culminates with the generation of actionable solutions, ready to be integrated into later decision-making workflows.
5 Foundational principles for the development of industrial large knowledge models
@@
As illustrated in Figure 3, we propose the ”6S Principle” as a guideline for the future development of ILKMs. ”6S Principle” encompasses six key components: ”Specialized Domain Knowledge”, ”Scrutability”, ”Safety”, ”Scalability”, ”Sustainability”, and ”Systematization & Standardization”. The details of the purpose, challenges, and opportunities for each principle are briefly discussed in Figure 3. All of these principles are crucial for ILKM’s successful application in industrial settings, ensuring that ILKMs meet the specific needs and challenges of smart manufacturing environments.
@@ 6 Conclusion
This paper presents the ILKM framework offering a comprehensive blueprint for integrating advanced AI, LLMs, LKL, and ML technologies with specialized industrial knowledge into the realm of smart manufacturing. It shows a big potential for the ILKM to integrate with the next generation of industrial systems to achieve more intelligence, efficiency, and resilience.

### v2

@@ Abstract
The recent emergence of large language models (LLMs) shows the potential for artificial general intelligence, revealing new opportunities in industry 4.0 and smart manufacturing. However, a notable gap exists in applying these LLMs in industry, primarily due to their training on general knowledge rather than domain-specific knowledge. Such specialized domain knowledge is vital for effectively addressing the complex needs of industrial applications. To bridge this gap, this paper proposes an Industrial Large Knowledge Model (ILKM) framework emphasizing their potential to revolutionize the industry in smart manufacturing. In addition, ILKMs and LLMs are compared from eight perspectives. Finally, the “6S Principle” is proposed as the guideline for ILKM development, and several potential opportunities are highlighted for ILKM deployment in smart manufacturing.
@@ 1 Introduction
In the era of Industry 4.0 [6], a paradigm shift is unfolding in the manufacturing sector, driven by the advent of smart manufacturing practices. At the heart of this revolution lies the integration of many advanced techniques, such as Industrial Big Data Analytics [18], Industrial Artificial Intelligence (IAI) [8], Machine Learning (ML) and Deep Learning (DL) [11, 15], Cyber-Physical Systems (CPS) [7], Industrial Internet of Things (IoT) [13], etc. These techniques are all aimed at enhancing efficiency, productivity, and flexibility in manufacturing processes.
Recent advances in large language models (LLMs) [19, 2] have showcased extraordinary capabilities in natural language processing for understanding, interpreting, and generating human language. However, a gap exists in the application of these LLMs within smart manufacturing, primarily because LLMs are predominantly trained on general knowledge, not domain knowledge, which may not be entirely suitable for the specific and often complex needs of industrial applications. Therefore, there is an urgent need for the development of an advanced foundation model leveraging the powers of LLMs and domain knowledge to address complex challenges in smart manufacturing.
Recognizing this gap, we propose an industrial large knowledge model (ILKM) framework for domain-driven data-centric industrial systems in smart manufacturing, and the “6S Principle” is proposed as the guideline for the development of ILKM. The role of ILKMs and the comparison with LLMs are discussed in detail. Through this exploration, we aim to provide a comprehensive understanding of the transformative power of ILKMs in the modern manufacturing landscape, highlighting their significance and opportunities as a cornerstone of the ongoing industrial revolution.
@@ 2 The role of industrial large knowledge models in smart manufacturing
In smart manufacturing, the deployment of ILKMs emerges as a pivotal element. Figure 1 shows the general process of how ILKM works within Industry 4.0, where ILKM functions at the core of this advanced manufacturing paradigm. The process begins with the acquisition and aggregation of a vast array of industrial data, derived from diverse industrial products. This data is categorized into two primary forms: human-interpretable data and structured machine-generated data. Leveraging technologies such as LLMs, a comprehensive Large Knowledge Library (LKL), along with various ML and Artificial Intelligence (AI) techniques, the ILKM serves as Artificial General Intelligence (AGI), which plays a vital role in enabling advanced and sophisticated data analytics and problem-solving. This, therefore, paves the way for a more insightful and informed decision-making process. Beyond this, the ILKM can also interface with and improve supply chain management, leading to more efficient, resilient, and customer-focused operations. Additionally, the solutions generated by the ILKM undergo evaluation by Subject Matter Experts (SMEs), who play a crucial role in validating and refining the relevant solutions, thereby aiding in the continual optimization of ILKM’s output. This iterative process of assessment and feedback is integral to ensuring the relevance and effectiveness of ILKM’s solutions. Overall, the ILKM underscores the transformative potential of data-driven approaches, offering detailed and comprehensive optimization, and enhancement directions for industrial products in smart manufacturing.
Figure 2: Industrial Large Knowledge Model Framework.
3 Industrial large knowledge model framework
The proposed ILKM framework, shown in Figure 2, provides a step-by-step guideline for developing and deploying ILKMs using industrial data for enhancing manufacturing capabilities in areas such as predictive maintenance, process optimization, quality control, engineering design, QA platform, data analytics, etc. The framework of the ILKM consists of four pivotal steps: (1) the construction of a LKL categorized by human-interpretable and structured machine-generated data; (2) the preparation of domain-specific instruction data; (3) the development of domain knowledge large language model based on the domain-specific data and domain instruction data; (4) the establishment of an intelligent domain expert machine learning system. As illustrated in Figure 2, the details about the ILKM framework are outlined as follows:
3.1 Large knowledge library construction
The initial step of constructing an ILKM involves the creation of an LKL. This library is pivotal for accommodating the breadth and diversity of industrial data, thus serving as a foundational resource for subsequent analytical tasks. During this phase, it’s essential to systematically categorize the data into domain-specific categories. Such organization enables researchers and data scientists to streamline their efforts, allowing for efficient retrieval of domain-specific data to inform the development of ML models tailored to address distinct industry-related challenges. Within these categories, based on the usage and nature of the data, industrial data can be divided into two primary types: human-interpretable and structured machine-generated. On the one hand, human-interpretable data comprises elements such as text documents, annotated images, coding scripts crafted by programmers, multimedia content, etc, which are inherently designed for human cognition. These data can be seen as insightful information or knowledge and it can be used for later domain knowledge machine learning model development. On the other hand, structured machine-generated data, comprising sensor readings, machine logs, operational parameters, etc, can be leveraged for analytical purposes in technical and industrial contexts.
3.2 Domain instruction data preparation
In the second step of the ILKM framework, the focus shifts to transforming domain-specific data (human-interpretable data from LKL) into structured domain instruction sets by achieving multi-modal data fusion and generating domain-centric knowledge. This transformation is crucial for enhancing the performance of LLMs in targeted domains. These structured instructions, vital for fine-tuning the LLM and retrieving domain knowledge, ensure the model is proficient in addressing domain-specific challenges and enhancing problem-solving capabilities. As depicted in Figure 2, the domain instruction data is organized into three parts: firstly, the domain instruction, which identifies the problem’s domain and may include sub-tags for refined categorization; secondly, the input, which clearly outlines the current problem; and thirdly, the output, which presents the corresponding solution.
3.3 Domain knowledge large language model development
The third step of the ILKM framework entails an initial pre-training of the base LLM with domain-specific data sourced from LKL. This pre-training imbues the LLM with rich domain-specific knowledge. Following this, the pre-trained LLM undergoes a fine-tuning process, guided by domain instructions, transforming it into a Domain Knowledge LLM proficient in the designated field. To refine the LLM’s expertise, some common techniques to enhance and train LLMs can be summarized as follows: reinforcement learning from human feedback (RLHF) [4], instruction tuning [9, 16], mixture of experts (MOE) [12], prompt engineering [1, 10], retrieval-augmented generation (RAG) [5] and leveraging attention mechanism [14], etc. The objective of this step is to build a robust LLM that possesses extensive domain knowledge and expertise. This LLM can then guide the development of new ML models that are capable of addressing complex challenges and real industrial problems.
3.4 Intelligent domain expert machine learning system
Upon the successful training of the Domain Knowledge LLM, the fourth step utilizes it as a domain expert for subsequent specialized model development. In this step, domain instruction data serves as the prompt, propelling the LLM to address specific analytical problems. The Domain Knowledge LLM, acting on these instruction inputs, proposes targeted solutions. Additionally, human experts may interact and intervene, offering strategic guidance to refine the LLM’s outputs. These solutions are then transferred to a coding-focused LLM [3, 17], which incrementally develops code aligned with the Domain Knowledge LLM’s insights, thereby creating a new ML model for specific problems. In addition, the structured machine-generated data serves as the dataset for new ML model training and testing. Finally, this step culminates with the generation of actionable solutions, ready to be integrated into later decision-making workflows.
@@ 3.4 Intelligent domain expert machine learning system
4 Discussion
This section discusses the comparison between ILKMs and LLMs and introduces the ”6S Principle” as a guideline for future ILKM development. It also highlights several potential opportunities for ILKM deployment in smart manufacturing.
4.1 Comparison between ILKMs and LLMs
The main difference between the ILKMs and LLMs lies in their purpose and functionality. The ILKM is designed to handle specific industrial tasks, utilizing relevant structured industrial data and domain-specific knowledge to provide precise, expert-level solutions, while LLMs are more generalized, leveraging extensive training on diverse textual data to solve language-related tasks like text generation, conversation, language translation, etc. To better illustrate the characteristics of ILKMs, a detailed comparison between ILKM and LLM is presented in Table 1. We compare and explain them from eight perspectives: data, purpose, applications, data privacy and security, domain-specific knowledge, integration and customization, scalability, and real-time decision-making.
@@ 4.1 Comparison between ILKMs and LLMs
4.2 Foundational principles for ILKM development
@@
As illustrated in Figure 3, we propose the ”6S Principle” as a guideline for the future development of ILKMs. ”6S Principle” encompasses six key components: ”Specialized Domain Knowledge”, ”Scrutability”, ”Safety”, ”Scalability”, ”Sustainability”, and ”Systematization & Standardization”. The details of the purpose, challenges, and opportunities for each principle are briefly discussed in Figure 3. All of these principles are crucial for ILKM’s successful application in industrial settings, ensuring that ILKMs can address specific needs and challenges faced in smart manufacturing.
4.3 Prospective and opportunity
There are several opportunities for developing ILKMs in the future of smart manufacturing. In the field of material discovery and synthesis, ILKMs can analyze historical data from the literature to summarize design guidelines and principles as domain-specific instructions. These instructions can then guide the design of new experiments, ultimately facilitating the discovery of new materials. In the area of engineering design, ILKM can assimilate knowledge from multi-modalities (text, 2D images, 3D shapes, sound, etc) gathered from historical products, then provide possible optimization directions for the designers and engineers to enhance new product performance. In the realm of prognostics and health management, ILKMs can analyze data from historical failures and maintenance strategies to aid in the diagnosis and prognosis of complex industrial machine systems, ultimately enabling predictive maintenance and lifecycle management. Furthermore, intelligent QA platforms can be developed across various industrial sectors with the revolution of ILKM. Instead of manual information searches performed by humans, ILKM can automatically retrieve relevant information and generate responses, thus assisting employees with their queries.
@@ 5 Conclusion
This paper presents a unified ILKM framework offering a comprehensive blueprint for integrating advanced AI, LLMs, LKL, and ML technologies with specialized industrial knowledge into the realm of smart manufacturing. It shows a big potential for the ILKM to integrate with the next generation of industrial systems to achieve more intelligence, efficiency, and resilience.